OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 • x + 4 • 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10 WILL MEDAL ASAP!!!

OpenStudy (anonymous):

And fan

OpenStudy (benlindquist):

You want to have 12 + 28) = 4 • x + 4 • 7

OpenStudy (anonymous):

so how do u do it

OpenStudy (zarkon):

\[ 4 x + 4 \cdot 7=4(x+7)\] answer should be obvious now

OpenStudy (benlindquist):

Thats better

OpenStudy (anonymous):

so 4

OpenStudy (zarkon):

\[4(3+7)=4(x+7)\]

OpenStudy (pawanyadav):

So x=3 By comparison

OpenStudy (anonymous):

sorry im having trouble undertsanding so 3 would be it

OpenStudy (anonymous):

im just really bad at math